Each day Mandela, who is in the third grade, eats lunch at school. He only likes liver (L) and onions (N),

and these provide him a utility of;

U(L,N) = ln (LN)

Liver costs $4.00 per serving, onions cost $2.00 per serving, and Mandela’s mother gives him $10.00 to

spend on lunch.

a. Give the equation for Mandela’s budget line. If liver is on the vertical axis, what is the

slope of the budget line?

b. Graph Mandela’s budget line. Place the number of liver servings on the vertical axis and

the number of onion servings on the horizontal axis.

c. On the same graph, draw Mandela’s indifference curve (free hand curve). Show the

concept of tangency and discuss your graph


1
Expert's answer
2021-04-23T12:08:14-0400

Solution:

a.). The equation of the budget line: M = PxX + PyY

M = Income

PxX = Price of onions = 2

PyY = Price of liver = 4

Budget line:

10 = 2X + 4Y

X-intercept = "\frac{M}{Px}" = "\frac{10}{4}" = 5 = (5, 0)


Y-intercept ="\frac{M}{Py} = \frac{10}{4 } = 2.5" = (0, 2.5)

 

The slope of the budget line = "\frac{\triangle Y}{\triangle X} = \frac{(0 - 2.5)}{(5 - 0)} = \frac{-2.5}{5} = - 0.5"

 

b.). Mandela’s budget line graph:






c.). The concept of tangency refers to a point where the indifference curve is tangent to the budget line. Tangent point is where the budget line is convex to indifference curve and it is where utility is maximum. It is a point which shows the best affordable combination of goods given the budget constraint.

 

The optimal solution is the point where the indifference curve cuts the budget line, which is (4, 0.5).